OpenStudy (anonymous):

Write a radical expression that has 4 terms (without variables), that are not in simplest form. Create this expression so that, once the terms are in simplest form, they can be combined to make an expression with one or two terms. Then, show all your work to

OpenStudy (cwrw238):

7sqrt48 + 5sqrt45 - 2sqrt48 - 2 sqrt45

OpenStudy (cwrw238):

simplifying: 7 sqrt(3*16) + 5sqrt((5*9) - 2 sqrt(3*16) - 2 sqrt(5*9) = 7sqrt3sqrt16 + 5sqrt5sqrt9 -2sqrt3sqrt16 - 2 sqrt5sqrt9 = 7*4 sqrt3 + 5*3 sqrt5 - 2*4sqrt3 - 2*3sqrt5 = 20sqrt3 + 9sqrt5